OpenStudy (anonymous):

4 1/4 + 2 3/7 can u please show me the steps? thanks

OpenStudy (anonymous):

4 1/4 + 2 3/7 4 1/4 = 17/4 2 3/7 = 17/7 17/4 + 17/7 = you need a common denominator (28 in this case) 119/28 + 68/28 = 187/28 = 6 19/28

OpenStudy (anonymous):

sorry but i wrote it wrong.. it is 4 4/14 + 2 3/7..

OpenStudy (anonymous):

I have 60/14 + 34/14 = 94/14.. and then what?

OpenStudy (anonymous):

4 4/14 = 60/14 2 3/7 = 17/7 60/14 + 34/14 = 94/14 you can divide both the numerator and denominator by 2 to get 47/7 and then 7 goes into 47 6 times with 5 left over so the final result is 6 5/7

OpenStudy (anonymous):

thanks